OpenStudy (timaashorty):

PLEASE HELP ME IF YOU CAN.. simplify the expression. -6+i/-5+i

OpenStudy (anonymous):

@tima u k this is spanish right

OpenStudy (timaashorty):

THESE ARE THE CHOICE ANSWERS: \[ A. \frac{ 31+i }{ 26 }\] \[ B. \frac{ 29+i }{ 26 }\] \[ C. \frac{ 31+i }{ 24 }\] \[ D. \frac{ 31+11i }{ 26 }\]

OpenStudy (timaashorty):

Oh I Thought i was on math ! lol